I am forgetting how to work this!
Evaluate the function f (x)=5x-7
f (2)=
f (0)=
f (-4)=
Help is much appreciated

Answer by fractalier(6550)

You merely substitute the value in the parentheses wherever you find an x and evaluate, so that for
f(x)=5x-7
we have
f(2) = 5(2) - 7 = 10 - 7 = 3
f(0) = 5(0) - 7 = 0 - 7 = -7
f(-4) = 5(-4) - 7 = -20 - 7 = -27